OpenStudy (nikki_skellington):

Convert 11.42424242 … to a rational expression in the form of a over b, where b ≠ 0. 11 and 42 over 999 11 and 42 over 9999 11 and 42 over 99 11 and 99 over 42

OpenStudy (phi):

I would ignore the 11 (for the time being) let x= 0.424242... and 100x= 42.42424...

OpenStudy (phi):

notice that 100x - x = 42.4242... -0.4242.. = 42 or 99x= 42 x= 42/99 so you have the fraction \[ 11 \ \frac{42}{99} \]

OpenStudy (nikki_skellington):

I get it now, thx so much!